1. Determine the equation of the line that passes through the points (-4, 3) and (2,2).
2+-4 = 2
2+(-3)= 5
y+y=2/5 {x+(-4)}
y+y=2/5x+8/5
y=2/5x+8/5-2/5=2/5x=5/5=1

2. Find the x and y intercepts and the slope of the line for 4x-3y=-6
-3y=-4x-6
-3y -4x =-6
___ ___ ___
-3 -3 3-
y=3/4x -2
=4/3 y=0,4
x=3,0


3. Graph y = -3/4x -2


4. Graph 3x+2y= -6

Answer by stanbon(75887)

. Determine the equation of the line that passes through the points (-4, 3) and (2,2).
You need the slope and y-intercept:
Slope = [3-2]/[-4-2]= -1/6
y=mx+b
2=(-1/6)2+b
2=-1/3 + b
7/3 = b
EQUATION:
y=(-1/6)x+(7/3)
-------------
2. Find the x and y intercepts and the slope of the line for 4x-3y=-6
3y=4x+6
y=(4/3)x+2
Slope = (4/3)
y-intercept = 2
--------------
3. Graph y = (-3/4)x -2
Let x=0, then y=-2 giving you the point (0,-2)
Let x=4, then y=-5 giving you the point (4,-5)
Plot those two points and draw a line thru them.
That is the graph.
---------------

4. Graph 3x+2y= -6
Let x=0 then y=-3: Plot the point (0,-3)
Let y=0, then x=-2; Plot the point (-2,0)
Draw a line thru those points.
That is the graph.
Cheers,
stan H.